OpenStudy (anonymous):

The difference of 3 times a number and -1/2 is -5/2. What is the number and -1/2 means -one half and -5/2 means - five over 2

OpenStudy (anonymous):

3a- -1/2= -5/2 is the problem

OpenStudy (anonymous):

Then isolate the a

OpenStudy (anonymous):

add 1/2 to -5/2

OpenStudy (anonymous):

you get 3a= 1/2 you can isolate a....... hope this helped!

OpenStudy (anonymous):

oops the 1/2 is negatve by the way

OpenStudy (anonymous):

so what would be the answer?

OpenStudy (anonymous):

no you have to do 3a= -1/2 divide both sides by 3 so a= 3 divided by -1/2

OpenStudy (anonymous):

ok Thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Does anyone have a different answer? I just want to make sure this is correct

OpenStudy (anonymous):

Question will be closed in 5 minutes

OpenStudy (anonymous):

by the way 3 divided by a -1/2 is 6 a=6